# Large Language Models Do Multi-Label Classification Differently

### v1

@@ Abstract
Multi-label classification is prevalent in real-world settings, but the behavior of Large Language Models (LLMs) in this setting is understudied. We investigate how autoregressive LLMs perform multi-label classification, with a focus on subjective tasks, by analyzing the output distributions of the models in each generation step. We find that their predictive behavior reflects the multiple steps in the underlying language modeling required to generate all relevant labels as they tend to suppress all but one label at each step. We further observe that as model scale increases, their token distributions exhibit lower entropy, yet the internal ranking of the labels improves. Finetuning methods such as supervised finetuning and reinforcement learning amplify this phenomenon. To further study this issue, we introduce the task of distribution alignment for multi-label settings: aligning LLM-derived label distributions with empirical distributions estimated from annotator responses in subjective tasks. We propose both zero-shot and supervised methods which improve both alignment and predictive performance over existing approaches. Code available at https://github.com/gchochla/LLM-multilabel-differently.
@@
Marcus Mathanks: equal contribution, Georgios Chochlakis11footnotemark: 1, Niyantha Maruthu Pandiyan, Jesse Thomason, Shrikanth Narayanan University of Southern California Correspondence: {mjma, chochlak}@usc.edu
@@ 1 Introduction
A key reason may be the incompatibility between the language modeling objective and the multi-label setting. LLMs are trained to generate probability distributions over vocabulary tokens via softmax normalization, naturally lending themselves to single-label settings by restricting the normalization to label tokens. In contrast, multi-label classification does not require label probabilities to sum to one. Instead, each label’s confidence can, in principle, be modeled independently, which prevalent LLMs are not trained to do, as their logits are meaningful only in relation to each other.
@@ 2.1 LLM Usage for Multi-label Predictions
[…]
@@ 2.2 Subjective Language Tasks
Figure 2: Top probabilities at each generation step when the last or an intermediate label is generated. Patterns are identical between the two settings, and bigger or finetuned models have clusters closer to 100%.
@@ 3 Datasets
We present both objective and subjective multi-label datasets. We use 10-shot prompts with Llama3 Dubey et al. (2024) (more details in §A). We apply softmax over initial label tokens to derive label probabilities at each step.
@@ 4 Multi-Label Mechanisms of LLMs
Figure 3: Second-highest probabilities at each generation step when the last or an intermediate label is generated. We also show the probability at the current step of the label that is actually predicted in the next step ($r$+1 pred), the probability at the next generation step of the second highest probability of the current step (intermediate @ $r$+1), and the percentage of cases the second-highest probability label at step $r$ and the prediction at $r$+1 is the same. LLM distributions show poor relative ranking, and little distinction between the last and intermediate settings.
We evaluate whether LLMs produce diverse, consistent, and informative probability distributions. Specifically, we investigate whether the predicted probabilities at each generation step reflect the relative confidence of the LLM and whether the relative ordering of labels provides insight into future predictions. To this end, we analyze the distribution of the top two predicted probabilities at each label generation step, along with the entropy of the distribution, allowing us to assess how spiky the distributions are, that is, how close the top probability is to 1 and how low the entropy is.
We also compare the top probabilities to evaluate whether their relative values reflect the model’s confidence. Crucially, we examine the second-highest probability and track how it evolves in the subsequent generation step. By distinguishing between steps where the model continues generating more labels (denoted as intermediate) and steps where it predicts the final label (denoted as last), we assess whether the second-highest probability provides a meaningful signal about the future behavior.
Finally, we test whether the relative order of the probabilities is informative by comparing the second-highest probability in the current generation step to that of the label generated in the next.
Figures 2 and 3 show the results based on the predicted probabilities for all datasets using Llama3 8B and 70B Base, Instruct, and with Supervised Finetuning Ouyang et al. (2022) (SFT; details in §A.5). We show only up to the second step to avoid clutter. Instances with only one shown generation step predicted only up to two labels. Corresponding entropy measures can be found in §D.2.
@@
We demonstrate that LLMs do not reliably pick the second highest label as their next prediction, even if they continue predicting. For instance, in SemEval, the label with the second highest probability in the first step is not predicted next between $48.1\%$ and $69.2\%$ of the time across models. In GoEmotions, this behavior occurs between $22.2\%$ and $49.8\%$ of cases. In fact, if we take the label with the second-highest probability in the current step $r$, and look at its probability in the next step $r$+1 (shown as intermediate @ $r$+1), we see that it is clusters at 0. Similarly, when we look at the probability of the label predicted in step $r$+1, and see how its probability looked in the previous step $r$ (shown as $r$+1 pred), its probability tends to also be clustered around 0. Notably, we find that if the second highest label at any step is not predicted as the next generated label, it will not be not predicted at all most of the time (see §D.3). While is in some sense expected, since each generated label is newly conditioned on the previously generated labels (we verify this in §D.6 by looking at the attention weights), it means that each generation step is only informative of the current label, since the relative ordering of predicted labels is not predictive of subsequent behavior.
@@
Finally, we report that the label type of in-context prompts greatly influences the rate of multi-label output. We show in Figure 6 how the percentage of multi-label (as opposed to single or no label) examples roughly corresponds to the percentage of multi-label output across models and datasets. Learning to predict multi-label outputs must be highlighted very clearly in the in-context examples, suggesting that single-label formats have dominated the training of the model. Overall, these analyses demonstrate that LLMs do not create well-calibrated distributions when generating multiple labels; instead, they generate spiky distributions, classifying labels one at a time.
@@ 5 Multi-Label Distribution Alignment
[…]
Model GoEmotions MFRC SemEval Perf Gold Pred Pred 2+ Perf Gold Pred Pred 2+ Perf Gold Pred Pred 2+ 8B Base 0.282 0.463 0.795 0.197 0.115 0.246 0.637 0.211 0.522 0.590 0.744 0.647 8B Instruct 0.434 0.520 0.903 0.217 0.285 0.380 0.877 0.074 0.537 0.624 0.779 0.675 70B Base 0.357 0.519 0.876 0.0 0.228 0.368 0.802 0.043 0.603 0.616 0.812 0.706 70B Instruct 0.396 0.554 0.856 0.440 0.314 0.420 0.819 0.550 0.588 0.627 0.844 0.758
Table 2: Micro F1 $\uparrow$ of linear probes trained and evaluated on gold labels (Gold), trained and evaluated on model predictions (Pred), and evaluated on predictions beyond the first generated label (Pred 2+). For comparison, we also show the performance of the model (Perf). Embeddings are from the last layer for the first generated label.
@@ 5.2.2 Test-Time Methods
Given model’s good calibration on binary tasks Niculescu-Mizil and Caruana (2005), we propose:
@@ 5.4 Results
The linear probing method ranks as the second best baseline, so the hidden states during first-label generation alone seem, at first glance, to contain enough information to perform well on the tasks. However, in Table 2, we present a more detailed analysis with linear probes. In addition to model and probing performance, we present the probes’ capability of predicting the predictions of the model themselves (i.e., the probes are trained on the predictions). We present the performance on the predictions on the Pred column, showing, as expected, much higher performance. However, when we look at how well the probes can predict any label after the first (Pred 2+), we see a substantial degradation in performance. Note that the task in theory becomes easier as we remove a label from the problem. This degradation suggests that linear probing performs well mostly due to its high accuracy of the first label and has little predictive power for any future labels, which aligns with our findings that LLMs predict labels one at a time. Even after supervised training, embeddings of the first label generation do not contain enough information to predict any subsequent labels.
@@ A.2 Multi-Label Datasets
The seven emotion “clusters” are: admiration, anger, fear, joy, optimism, sadness, and surprise.
@@ D.1 Probabilities: Alternative view
Figure 8: Top two probabilities at each generation step $r$ (up to two for brevity) when the last label is generated, or when a middle label is generated. Shown are for four datasets, one per row. In each row, the bottom subfigure shows the top probability, and the top the second highest probability, in addition to the probability of the label that was actually predicted next at the current step ($r$+1 pred), and the probability at the next generation step of the second highest probability (mid @ $r$+1). Also shown is the percentage of cases the second-highest probability label at $r$ and the prediction at $r$+1 were the same. Instances with only one shown generation step predicted only up to two labels.
@@ D.2 Entropy of Predictions
Figure 9: Entropies of prediction distributions at each generation step $r$ when the last label is generated, or when a middle label is generated.
@@ D.3 Inconsistencies in second highest label scores
In Figure 10, we study in more detail the consistency of the second-highest probability label, excluding the instances where it was not predicted at all, and show the histograms for each generation step. We find that increasing the model size improves the rate at which that label is predicted right after it is ranked second, as Llama3 70B Instruct predicts the label with the second-highest probability as the second label $65$% of the time compared to approximately $50\%$ of the time with 8B Instruct. This indicates that with scale, the relative ordering of labels improves.
@@ D.4 Alignment of Llama3 8B
[…]
@@ D.6 Attention to Input vs Labels
We present the average attention to tokens in the prompt for models, when they generate the second or higher label. We intend to examine how much the models attend to the previous labels generated, establishing empirically the intuition that because of language modeling, the answers of the model deviate from whatever can be gauged from the first generated label token distribution. Table 6 shows that, on average, an order of magnitude higher weights are found in the label part of the prompt compared to the input (which also includes other labels because of the demonstrations). Attending to the format of the response is a plausible confounder, so we also check the attention specifically to the first label tokens. This suggests that, indeed, subsequent labels are conditioned on the previous generations. We note that even though average attention is lower on the input, cumulative attention is still greater, with approximately a $80\%$/$20\%$ split in favor of the input, which is usually an order of magnitude or more longer than the labels themselves, again suggesting that a lot of attention weights are accumulated on the generated labels.

### v2

@@ Abstract
Multi-label classification is prevalent in real-world settings, but the behavior of Large Language Models (LLMs) in this setting is understudied. We investigate how autoregressive LLMs perform multi-label classification, focusing on subjective tasks, by analyzing the output distributions of the models at each label generation step. We find that the initial probability distribution for the first label often does not reflect the eventual final output, even in terms of relative order and find LLMs tend to suppress all but one label at each generation step. We further observe that as model scale increases, their token distributions exhibit lower entropy and higher single-label confidence, but the internal relative ranking of the labels improves. Finetuning methods such as supervised finetuning and reinforcement learning amplify this phenomenon. We introduce the task of distribution alignment for multi-label settings: aligning LLM-derived label distributions with empirical distributions estimated from annotator responses in subjective tasks. We propose both zero-shot and supervised methods which improve both alignment and predictive performance over existing approaches. We find one method – taking the max probability over all label generation distributions instead of just using the initial probability distribution – improves both distribution alignment and overall F1 classification without adding any additional computation.
@@
Marcus Mathanks: Equal contribution. Code is available at https://github.com/gchochla/LLM-multilabel-differently., Georgios Chochlakis11footnotemark: 1, Niyantha Maruthu Pandiyan, Jesse Thomason, Shrikanth Narayanan University of Southern California Correspondence: {mjma, chochlak}@usc.edu
@@ 1 Introduction
A key reason may be the incompatibility between the language modeling objective and the multi-label setting. LLMs are trained to generate probability distributions over vocabulary tokens via softmax normalization for the immediate next token, naturally lending themselves to single-label settings, such as by restricting the normalization to label tokens. In contrast, multi-label classification does not require label probabilities to sum to one. Instead, each label’s confidence can, in principle, be modeled independently. This runs counter to how LLMs are trained, as their logits are meaningful only in relation to each other.
@@
  • We highlight the zero-shot approach of max-over-generations, which improves both distribution alignment and F1 classification for no additional computation. This method involves setting a label’s probability to its max value across all label generations rather than its value in a single label distribution.
@@ 2.1 LLM Usage for Multi-label Predictions
To the best of our knowledge, Niraula et al. (2024) is the only work to explicitly investigate LLM multi-label classification Chen et al. (2022) in niche domains. Bețianu et al. (2024) explored a multi-label framework for finetuning BERT and Jung et al. (2023) trained a classifier on top of T5 encodings directly for multi-label classification rather than relying on model text generation. The two well-studied forms of multi-label classification are extreme multi-label classification (XMLC;  Zhu and Zamani 2024), where models must assign many labels to a document from a very large label set (1000+ labels), and hierarchical multi-label classification Tabatabaei et al. (2025), where labels are subdivided into sub-labels recursively. Subjective multi-label classification is relatively unexplored Chochlakis et al. (2024). We thoroughly investigate LLMs in these settings by analyzing their classification patterns across datasets.
@@ 2.2 Subjective Language Tasks
Figure 2: Top probabilities at each generation step when the last or an intermediate label is generated. Patterns are identical between the two settings, and bigger or finetuned models have clusters closer to 100%. A single step only is shown when only up to labels were generated for all examples in a specific setting.
@@ 3 Datasets
We present both objective and subjective multi-label datasets. We use 10-shot prompts with Llama3 Dubey et al. (2024) (more details in §A). We apply softmax over initial label tokens to derive label probabilities at each step. It is well known that several different tokens can correspond to the same concept Holtzman et al. (2022), such as “happy”, “Happy”, and “ happy”, and found that selecting the highest logit score across all same-concept tokens as a given label’s logit value was the most effective way to capture model belief.
@@
Dataset Annotators (per example) Cohen’s Kappa 0 labels 1 label 2 labels 3+ labels GoEmotions 81 (3.58) 0.27 29% 62% 8% 1% MFRC 6 (2.99) 0.21 78% 18% 3% <<1% SemEval – (–) 1% 13% 40% 46%
Table 1: Annotation statistics and label distributions. The public release of SemEval does not include individual annotator labels, only aggregates.
@@ 4 Multi-Label Mechanisms of LLMs
Figure 3: Second-highest probabilities at each generation step when the last or an intermediate label is generated. We also show the probability at the current step of the label that is actually predicted in the next step (rr+1 pred), the probability at the next generation step of the second highest probability of the current step (intermediate @ rr+1), and the percentage of cases the second-highest probability label at step $r$ and the prediction at rr+1 is the same. LLM distributions show poor relative ranking, and little distinction between the last and intermediate settings. A single step only is shown when only up to labels were generated for all examples in a specific setting.
We evaluate whether LLMs produce diverse, consistent, and informative probability distributions. Specifically, we investigate whether the predicted probabilities at each generation step reflect the relative confidence of the LLM and whether the relative ordering of labels provides insight into future predictions. To this end, we analyze the distribution of the top two predicted probabilities at each label generation step, along with the entropy of the distribution, allowing us to assess how spiky the distributions are, that is, how close the top probability is to 1 and how low the entropy is. That is, we take the output probabilities of the model at each generation step where a label starts being predicted (if the LLM breaks a label into multiple tokens, then we take into account the probabilities for only the first token), extract the top two probabilities for further analysis, and also compute the entropy of the entire distribution.
We also compare the top probabilities to evaluate whether their relative values reflect the model’s confidence. Crucially, we examine the second-highest probability and track how it evolves in the subsequent generation step, and importantly how often the corresponding label is predicted next, as would be expected. By distinguishing between steps where the model continues generating more labels (denoted as intermediate) and steps where it predicts the final label (denoted as last), we assess whether the second-highest probability provides a meaningful signal about future behavior.
Finally, we test whether the relative order of the probabilities is informative by comparing the second-highest probability in the current generation step to that of the label generated in the next. That is to say, we look at the next generation step, see the label that was actually predicted, and then compare that label’s probability in the current generation step compared to the probability of the second-highest label in the current step.
Figures 2 and 3 show the results based on the predicted probabilities for all datasets using Llama3 8B and 70B Base, Instruct, and with Supervised Finetuning Ouyang et al. (2022) (SFT; details in §A.5). We show only up to the second step to avoid clutter. Corresponding entropy measures can be found in §D.2. We highlight key findings below.
@@
We demonstrate that LLMs do not reliably pick the second highest label as their next prediction, even if they continue predicting. For instance, in SemEval, the label with the second highest probability in the first step is not predicted next between $48.1\%$ and $69.2\%$ of the time across models. In GoEmotions, this behavior occurs between $22.2\%$ and $49.8\%$ of cases. In fact, if we take the label with the second-highest probability in the current step $r$, and look at its probability in the next step rr+1 (shown as intermediate @ rr+1), we see that it is clusters at 0. Similarly, when we look at the probability of the label predicted in step rr+1, and see how its probability looked in the previous step $r$ (shown as rr+1 pred), its probability tends to also be clustered around 0. Notably, we find that if the second highest label at any step is not predicted as the next generated label, it will not be not predicted at all most of the time (see §D.3). While is in some sense expected, since each generated label is newly conditioned on the previously generated labels (we verify this in §D.6 by looking at the attention weights), it means that each generation step is only informative of the current label, since the relative ordering of predicted labels is not predictive of subsequent behavior.
@@
Finally, we report that the label type of in-context prompts greatly influences the rate of multi-label output. We show in Figure 6 how the percentage of multi-label (as opposed to single or no label) examples roughly corresponds to the percentage of multi-label output across models and datasets. Learning to predict multi-label outputs must be highlighted very clearly in the in-context examples, suggesting that single-label formats have dominated the training of the model. Overall, these analyses show that LLMs do not create well-calibrated distributions when generating multiple labels; instead, they generate spiky distributions, classifying labels one at a time.
@@
Generalizability
To ensure our findings generalize to other model families, we replicate the main results for the Qwen 2.5 Team (2024) family of models in §D.7, showing identical results to the Llama family. Moreover, we experimented with an LLM with multiple decoding heads, Medusa Cai et al. (2024). Given its ability to predict multiple tokens at a time, the aforementioned behaviors might not be present in such models. Contrary to this assumption, we show in §D.8 that the model behaves in identical ways. Finally, in §D.9 we examine whether the label order in the instructions has a role in these phenomena, finding strong effects.
@@ 5 Multi-Label Distribution Alignment
Table 2: Distribution alignment scores for Llama3 70B Instruct on single and multi-label datasets between LLM and human distributions. $\textbf{F1}\uparrow$ is the example-F1 score. N/A: Not applicable to single-label setting. N/S: Not supplied to avoid clutter, and due to environmental considerations, since single-label settings are not our focus.
Figure 7: Micro F1 $\uparrow$ of linear probes trained and evaluated on gold labels (Gold), trained and evaluated on model predictions (Pred), and evaluated on predictions beyond the first generated label (Pred 2+). For comparison, we also show the performance of the model (Perf). Embeddings are from the last layer for the first generated label.
@@ 5.2.2 Test-Time Methods
Findings from Niculescu-Mizil and Caruana (2005) indicate that binary tasks are generally well-calibrated. Even though modern LLMs are very different from the basic neural networks tested in this paper, we were inspired to design several different approaches that “break down” multi-label classification into smaller steps. For these methods, we investigated Monte Carlo sampling methods but found this approach simply added noise over directly calculating the label probabilities analytically.
@@ 5.4 Results
The linear probing method ranks as the second best baseline, so the hidden states during first-label generation alone seem, at first glance, to contain enough information to perform well on the tasks. However, in Figure 7, we present a more detailed analysis with linear probes. In addition to model and probing performance, we present the probes’ capability of predicting the predictions of the model themselves (i.e., the probes are trained on the predictions). We present the performance on the predictions on the Pred column, showing, as expected, much higher performance. However, when we look at how well the probes can predict any label after the first (Pred 2+), we see a substantial degradation in performance. Note that the task in theory becomes easier as we remove a label from the problem. This degradation suggests that linear probing performs well mostly due to its high accuracy of the first label and has less predictive power for any future labels, which aligns with our findings that LLMs predict labels one at a time. Even after supervised training, embeddings of the first label generation do not contain enough information to predict any subsequent labels.
@@ A.2 Multi-Label Datasets
The seven emotion “clusters” are: admiration (includes pride, gratitude, relief, approval, realization), anger (includes disgust, annoyance, disapproval), fear (includes nervousness), joy (includes amusement, excitement, love), optimism (includes desire, caring), sadness (includes remorse, embarrassment, disappointment, grief), and surprise (includes confusion, curiosity). The clustering was performed using the hierarchical clustering algorithm, applied on the correlations between emotions, as described in Demszky et al. (2020).
@@ D.1 Probabilities: Alternative view
Figure 9: Top two probabilities at each generation step $r$ (up to two for brevity) when the last label is generated, or when a intermediate label is generated. Shown are for four datasets, one per row. In each row, the bottom subfigure shows the top probability, and the top the second highest probability, in addition to the probability of the label that was actually predicted next at the current step (rr+1 pred), and the probability at the next generation step of the second highest probability (intermediate @ rr+1). Also shown is the percentage of cases the second-highest probability label at $r$ and the prediction at rr+1 were the same. A single step only is shown when only up to labels were generated for all examples in a specific setting.
@@ D.2 Entropy of Predictions
Figure 10: Entropies of prediction distributions at each generation step $r$ when the last label is generated, or when a intermediate label is generated, shown for the first two label generation steps. A single step only is shown when only up to labels were generated for all examples in a specific setting.
@@ D.3 Inconsistencies in second highest label scores
In Figure 11, we study in more detail the consistency of the second-highest probability label, excluding the instances where it was not predicted at all, and show the histograms for each generation step. We find that increasing the model size improves the rate at which that label is predicted right after it is ranked second, as Llama3 70B Instruct predicts the label with the second-highest probability as the second label 6565% of the time compared to approximately $50\%$ of the time with 8B Instruct. This indicates that with scale, the relative ordering of labels improves.
@@ D.4 Alignment of Llama3 8B
Table 5: Distribution alignment scores for Llama 3 8B on single and multi-label datasets between LLM and human distributions. $\textbf{F1}\uparrow$ is the example-F1 score. N/S: Not supplied to avoid clutter.
@@ D.6 Attention to Input vs Labels
We present the average attention to tokens in the prompt for models, when they generate the second or higher label. We intend to examine how much the models attend to the previous labels generated, establishing empirically the intuition that because of language modeling, the answers of the model deviate from whatever can be gauged from the first generated label token distribution. Table 6 shows that, on average, an order of magnitude higher weights are found in the label part of the prompt compared to the input (which also includes other labels because of the demonstrations). Attending to the format of the response is a plausible confounder, so we also check the attention specifically to the first label tokens. This suggests that, indeed, subsequent labels are conditioned on the previous generations. We note that even though average attention is lower on the input, cumulative attention is still greater, with approximately a $80\%$$20\%$ split in favor of the input, which is usually an order of magnitude or more longer than the labels themselves, again suggesting that a lot of attention weights are accumulated on the generated labels.
D.7 Results on Qwen
In this section, we replicate our main Llama findings for the Qwen 2.5 Team (2024) family, and in particular for:
  • Qwen 2.5 7B Base (Qwen/Qwen2.5-7B)
  • Qwen 2.5 7B Instruct (Qwen/Qwen2.5-7B-Instruct)
  • Qwen 2.5 72B Base (Qwen/Qwen2.5-72B)
  • Qwen 2.5 72B Instruct (Qwen/Qwen2.5-72B-Instruct)
We present our results for the top two probabilities at each step in Figures 13 and 14, and our linear probing results in Figure 15. We see identical with the Llama family, and Qwen can even be said to be more spiky.
Figure 13: Top probabilities at each generation step when the last or an intermediate label is generated. Patterns are identical between the two settings, and bigger or finetuned models have clusters closer to 100%. A single step only is shown when only up to labels were generated for all examples in a specific setting.
Figure 14: Second-highest probabilities at each generation step when the last or an intermediate label is generated. We also show the probability at the current step of the label that is actually predicted in the next step (rr+1 pred), the probability at the next generation step of the second highest probability of the current step (intermediate @ rr+1), and the percentage of cases the second-highest probability label at step $r$ and the prediction at rr+1 is the same. LLM distributions show poor relative ranking, and little distinction between the last and intermediate settings. A single step only is shown when only up to labels were generated for all examples in a specific setting.
Figure 15: Micro F1 $\uparrow$ of linear probes on Qwen 2.5 trained and evaluated on gold labels (Gold), trained and evaluated on model predictions (Pred), and evaluated on predictions beyond the first generated label (Pred 2+). For comparison, we also show the performance of the model (Perf). Embeddings are from the last layer for the first generated label.
D.8 Medusa: Multiple Decoding Heads
In this section, we present some results from a model with multiple decoding heads, Medusa (specifically FasterDecoding/medusa-1.0-zephyr-7b-beta). Shown in Figures 16, we see that this model shows behavior similar to Llama 3 and Qwen 2.5.
Figure 16: Top and second-highest probabilities at each generation step when the last or an intermediate label is generated. We also show the probability at the current step of the label that is actually predicted in the next step (rr+1 pred), the probability at the next generation step of the second highest probability of the current step (intermediate @ rr+1), and the percentage of cases the second-highest probability label at step $r$ and the prediction at rr+1 is the same. Patterns are identical between the two settings, and bigger or finetuned models have clusters closer to 100%. LLM distributions show poor relative ranking, and little distinction between the last and intermediate settings. A single step only is shown when only up to labels were generated for all examples in a specific setting.
D.9 Alphabetical Order
One potential confounding factor in the generation of many labels is their alphabetical order. By default, the labels are presented in an alphabetical order in the instructions and in the demonstrations, as any other mode of presentation would require justification. However, the strong alphabetical priors of the models coupled with the presentation in alphabetical order might be a strong driver of the phenomena we see. Therefore, in this section we present an analysis on how often that happens, as well as randomizing the order of the labels and examining whether the labels follow an alphabetical order or the new order of the instructions. Results are shown in Table 7, aggregated across Llama3 8B Base and Instruct. We see that proper alphabetical order ossifies the predictions of the model, but the reverse alphabetical order, which is also a regular patter, also does the same yet to a lesser extend. Future research can examine whether randomizing the prompt and aggregating across different orders might help to extract probabilities from the first logits, but this still requires multiple runs, making it more expensive that Max-over-Generations.
Setting Alphabetical (%) Prompt (%)
Alphabetical 96.4 -
Random 35.2 40.1
Reverse 15.9 71.9
Table 7: Percentage of predictions that follow alphabetical order and the order of the labels in the instructions in three settings: Alphabetical order of labels, Random order of labels (3 different seed) and Reverse alphabetical order of labels.